naiezire:

what's 8divided by 2(2+2)=

FexrlessRxby:

So the equation for this is 8/2(2+2) correct?

naiezire:

Yes it is solve it

Extrinix:

\(\dfrac{8}{2}(2+2)\) Easier way to show it.

FexrlessRxby:

I'd rather not, considering you just demanded for me to solve it.

naiezire:

I need sum metals so I'm earing them

naiezire:

Ok

supie:

first i would solve 8/2 which is the same thing as 8÷2 do you know what 8÷2 is?

naiezire:

Yes thats how I would solve it

IceDragon123:

oh uh so \[8 divided by 2\] that's easy, right?

IceDragon123:

ok so what does divided by mean do you know?

CripQUEZZ:

So what’s 8/2=?

CripQUEZZ:

@icedragon123 wrote:
ok so what does divided by mean do you know?
That’s a dumb question

naiezire:

Solve the problem

CripQUEZZ:

@naiezire wrote:
Solve the problem
We cant solve it for u

IceDragon123:

We can only help you get to it not say it 4 u

XxXNessalulbaddieXxX:

damn im wondering where the fraction came from

KiWiPoP1:

do inside the parenthesis first, and multiply the 2 on the outside and whatever you get from the "2+2" on the inside

naiezire:

4

CripQUEZZ:

@alwayscutegang wrote:
who u reported
They reported u bc ur giving direct answers

KiWiPoP1:

and what's 2x4

IceDragon123:

yeah

naiezire:

8

IceDragon123:

good

IceDragon123:

and 8 divided by 2

KiWiPoP1:

now all you need to do left is divide your original 88 by your new 8

CripQUEZZ:

@icedragon123 wrote:
and 8 divided by 2
No

KiWiPoP1:

@kiwipop1 wrote:
now all you need to do left is divide your original 88 by your new 8
*8

IceDragon123:

then why it say that on the question?

KiWiPoP1:

I don't think you're doing the math right if you're getting 2

CripQUEZZ:

@icedragon123 wrote:
then why it say that on the question?
U multiply by 2 not divide by it

naiezire:

Exactly

KiWiPoP1:

so your final equation would be 8/8

CripQUEZZ:

Yeah

KiWiPoP1:

you can rlly just put 2(2+2) into a calc, and boom

CripQUEZZ:

@icedragon123 wrote:
how the heck did u guess
There’s only two genders it’s not hard

KiWiPoP1:

yeah its pretty much solved anyways

IceDragon123:

let s solve the problem then

CripQUEZZ:

@icedragon123 wrote:
let s solve the problem then
We alr did

KiWiPoP1:

it's already solved

supie:

kiwipop1 wrote:
so your final equation would be 8/8
this is what u got as the answer....??

KiWiPoP1:

@supie wrote:
kiwipop1 wrote:
so your final equation would be 8/8
this is what u got as the answer....??
well I wasn't going to just say the answer, that's for them to figure out ykyk

supie:

how'd you get 8/8 tho? yk the question is \(\frac82(2+2)\) right?

KiWiPoP1:

I think it was the typa thing like 8x2(2+2)

CripQUEZZ:

@supie wrote:
how'd you get 8/8 tho? yk the question is \(\frac82(2+2)\) right?
I don’t think it’s a fraction it doesn’t say it is

supie:

8÷2=8/2

supie:

so \(8÷2(2+2)\) is the same thing as \(\frac82(2+2)\)

KiWiPoP1:

so then you'd still get the same answer either way I take it

CripQUEZZ:

Yeah

supie:

kiwipop1 wrote:
so then you'd still get the same answer either way I take it
but im confused on what you did what did you get for your answer ?

KiWiPoP1:

I got 1 for my answer

supie:

how lol

CripQUEZZ:

I thought it was 16

KiWiPoP1:

2(2+2)= 8 8/8=1

supie:

@cripquezz wrote:
I thought it was 16
same

supie:

@kiwipop1 wrote:
2(2+2)= 8 8/8=1
but it's 8/2(2+2) not 2(2+2)

supie:

this is what i did: \(\color{aqua}{\frac82}\color{tomato}{(}\color{lime}{2+2}\color{tomato}{)}\) first, solve 2+2 which equals 4 \(\color{aqua}{\frac82}\color{tomato}{(}\color{lime}{4}\color{tomato}{)}\) solve 8/2 which is the same thing as 8÷2 and 8÷2=4 \(\color{aqua}{4}\color{tomato}{(}\color{lime}{4}\color{tomato}{)}\) all that's left is \(\color{aqua}{4}\times\color{lime}{4}\), and \(4\times4=16\) so the answer would be 16

CripQUEZZ:

Bc if u distribute 8/2 into (2 + 2) it equals 16/2 and then u simplify and add and get 16 right?

CripQUEZZ:

@cripquezz wrote:
Bc if u distribute 8/2 into (2 + 2) it equals 16/2 and then u simplify and add and get 16 right?
Well (16/2 + 16/2)

KiWiPoP1:

I guess it depends on how @naiezire meant to put it as, if it was meant in the way you put it then I'm just wrong

supie:

if you used the distributive property then yeah

KiWiPoP1:

I thought the 8 was it's own thing and the 2(2+2) was also its own thing

supie:

kiwipop1 wrote:
I guess it depends on how naiezire meant to put it as, if it was meant in the way you put it then I'm just wrong
the question says "what's 8divided by 2(2+2)=" no?

supie:

@kiwipop1 wrote:
I thought the 8 was it's own thing and the 2(2+2) was also its own thing
oh, well it's 8/2(2+2)

KiWiPoP1:

yeah but if that was meant for your way, why isn't it "what is 8/2 (2+2)" or smth like that, the way he put it made it seem that 2(2+2) was together

jhonyy9:

why is this so difficile to solve it ? what say the text of this problem ? ,,what's 8 divided by 2(2+2) ?" 2+2=4 2×4=8 8 divided by 8 = 1 that's ALL !!!

KiWiPoP1:

LMAO ok so I got it, nice nice nice 😎

KiWiPoP1:

well eh